OpenStudy (anonymous):

the half life of a certain substance is 16 years. How long will it take for a sample of this substance to decay to 84% of its original amount? Use the exponential decay model THANKS!!!

OpenStudy (anonymous):

you can solve \[\huge \left(\frac{1}{2}\right)^{\frac{t}{16}}=0.84\]to get your answer

OpenStudy (anonymous):

you got that?

OpenStudy (photon336):

do you know how to solve this?